### v2

# OmniNWM: Omniscient Driving Navigation World Models

Abstract
Autonomous driving world models are expected to work effectively across three core dimensions: state, action, and reward. Existing models, however, are typically restricted to limited state modalities, short video sequences, imprecise action control, and a lack of reward awareness. In this paper, we introduce OmniNWM, an omniscient panoramic navigation world model that addresses all three dimensions within a unified framework. For state, OmniNWM jointly generates panoramic videos of RGB, semantics, metric depth, and 3D occupancy. A flexible forcing strategy enables high-quality long-horizon auto-regressive generation. For action, we introduce a normalized panoramic Plücker ray-map representation that encodes input trajectories into pixel-level signals, enabling highly precise and generalizable control over panoramic video generation. Regarding reward, we move beyond learning reward functions with external image-based models: instead, we leverage the generated 3D occupancy to directly define rule-based dense rewards for driving compliance and safety. Extensive experiments demonstrate that OmniNWM achieves state-of-the-art performance in video generation, control accuracy, and long-horizon stability, while providing a reliable closed-loop evaluation framework through occupancy-grounded rewards.
Figure 1: Versatile capabilities of OmniNWM. (a) State: Given a conditional image and ego trajectory, OmniNWM jointly generates comprehensive multi-modal outputs, including panoramic RGB, semantic, metric depth, and 3D occupancy videos. (b) Action: Conditioned on different input trajectories, OmniNWM facilitates precise panoramic camera control by converting them into normalized Plücker ray-maps as pixel-level guidance. (c) Reward: OmniNWM enables unbounded long-term navigation through a closed-loop pipeline: the planning trajectory guides the multi-modal generation, while dense rewards are derived from the generated 3D semantic occupancy. Note that all the input trajectories are encoded into normalized Plücker ray-maps before guiding the generation process, even when not explicitly visualized.
1 Introduction
Recent advancements in world models have demonstrated significant potential for autonomous driving, facilitating high-fidelity simulation of complex environments and controllable navigation for autonomous agents [72, 5, 43, 99, 73, 23]. These models typically predict plausible future environmental states conditioned on historical observations and actions, then evaluate action feasibility via reward functions [17, 18, 73, 43, 53].
Method State Action Reward
Modalities Length
GenAD [106] R 6 Waypoint
DrivingGPT [10] R 60 Waypoint
Drivedreamer [72] R 32 Velocity,Angle
Drivedreamer-2 [102] R 26 Velocity,Angle
Panacea [78] R 8 Waypoint
MagicDrive [16] R 60 Waypoint
Dist-4D [21] R,D 17 Waypoint
WoVoGen [51] R,O 6 Volumetric-based
OccScene [44] R,O 8 Volumetric-based
UniScene [43] R,L,O 24 Volumetric-based
Drive-WM [73] R 8 Waypoint External Image-based
OmniNWM (Ours) R,S,D,O 321 Normalized Plücker Ray-map-based Integrated Occupancy-based
Table 1: Comparison of world model capabilities. We compare OmniNWM with existing approaches across state, action, and reward dimensions. ‘R’, ‘S’, ‘D’, ‘L’, ‘O’ denote RGB, semantic, depth, LiDAR, and occupancy modalities, respectively.
Nevertheless, building robust and generalizable driving world models remains challenging due to the complexity of real-world 3D environments [22, 55, 91, 110, 81, 90, 52, 43]. Specifically, as shown in Table 1, existing methods face three core limitations: 1) State: Driving world models must integrate diverse data formats over sufficient temporal frames for closed-loop evaluation and downstream tasks [43, 16, 21]. Current approaches predominantly rely on single-modality RGB videos with limited length, which fail to capture the geometric and semantic complexity of large-scale 3D scenes [17, 18, 16, 73, 53]. 2) Action: Precise control for generating panoramic videos remains under-explored. Existing approaches often employ sparse action encodings (e.g., trajectory waypoints) or heavy volumetric conditions that hinder long-term consistency and precise viewpoint manipulation  [72, 102, 16, 73]. 3) Reward: Effective reward functions are essential for evaluating navigation performance, yet few methods have integrated them into a unified framework [72, 17, 102, 78, 43]. Although some recent studies propose image-based reward functions, they often rely on external models and fail to provide precise evaluations in driving environments [73]. These limitations collectively limit the robustness and generalization of current methods.
To address these challenges, we propose OmniNWM, an omniscient navigation world model for autonomous driving. OmniNWM unifies multi-modal forecasting over unbounded long-term sequences, precise and generalizable panoramic camera control, and seamless integration with 3D occupancy-grounded dense driving rewards within a single framework. As illustrated in Figure 1 (a), OmniNWM jointly generates panoramic videos of RGB, semantic, metric depth, and 3D semantic occupancy in an autonomous driving scenario. A flexible forcing strategy injects independent multi-level noise during training, enabling robust long-term generation beyond ground-truth sequence lengths (i.e., 321 vs. 241 frames in Figure 1(c)) and supporting frame- or clip-level auto-regressive inference. Figure 1 (b) illustrates the precise action control for panoramic video generation. In contrast to the sparse representation of trajectory waypoints used in prior work [72, 102, 18], we encode action sequences into normalized panoramic Plücker ray-maps—a unified pixel-level representation that facilitates accurate and generalizable control. Moreover, this normalized panoramic Plücker encoder is lightweight and introduces minimal overhead. Finally, as illustrated in Figure 1(c), OmniNWM supports closed-loop navigation by leveraging its generated 3D occupancy to provide dense waypoint-wise rewards, while planned trajectories iteratively guide the generation of future multi-modal predictions. Our contributions are summarized as follows:
  • We propose OmniNWM, an omniscient navigation world model that supports comprehensive forecasting, precise action control, and occupancy-grounded dense rewards within a unified framework.
  • A normalized panoramic ray-map representation is introduced to enable precise and generalizable camera control.
  • Driven by a flexible forcing strategy, OmniNWM achieves long-term auto-regressive generation beyond GT length.
  • Leveraging jointly generated 3D occupancy, OmniNWM naturally supports closed-loop navigation and evaluation through integrated, occupancy-grounded dense rewards.
Experiments show that OmniNWM achieves state-of-the-art performance in generation quality and control precision. Furthermore, the model exhibits remarkable generalization across diverse and challenging settings, including panoramic generation conditioned on novel trajectories, unbounded generation beyond GT sequence lengths, and zero-shot generation across different datasets and camera configurations.
Figure 2: Architecture overview of OmniNWM. Our model takes as input a historical trajectory, a reference image, and flexible multi-level noise applied view-wise and frame-wise. It simultaneously forecasts panoramic RGB, semantic, and metric depth videos (Section 3.1.1), along with 3D occupancy (Section 3.1.2) and a future planning trajectory (Section 3.1.3). Note that input trajectories guide the generation of the current video clip, while future trajectories are forecasted for the next clip, forming a closed-loop system. The input trajectories with camera poses are encoded into normalized panoramic Plücker ray-maps, providing a unified pixel-level description and enabling precise action control (Section 3.2). A flexible forcing strategy injects independent multi-level noise at the view and frame levels to support robust and flexible auto-regressive generation (Section 3.3)
2 Related Work
World Models for Autonomous Driving. Recent progress in driving world models has advanced visual realism and controllability [94, 18, 40, 16, 42, 73, 8, 92, 84, 19, 93, 60, 103, 53, 43, 70]. Early efforts like DriveDreamer [72] leverage multi-stage pipelines to incorporate traffic rules. MagicDrive [17] introduces cross-view attention for street-view synthesis, while Vista [18] improves generation via latent replacement strategies. UniScene [43] proposes a hierarchical occupancy-centric framework. Other notable works include ReSim [87] for simulation data augmentation, Drive-WM [73] for multi-view forecasting, and GEM [23] for egocentric prediction. However, existing models typically lack unified multi-modal generation (e.g., semantic map, 3D occupancy)  [72, 17, 18] and integrated reward modeling for closed-loop evaluation [102, 16, 43]. Our work addresses these limitations with long-horizon, multi-modal generation, precise control, and integrated reward modeling.
Camera-controlled Video Generation. Camera control has become a critical technique for guiding generative models in dynamic 3D environments [24, 25, 75, 37, 4, 89, 29]. CameraCtrl [24] introduces a plug-and-play module to parameterize camera trajectories. MotionCtrl [75] decouples camera and object motion for independent control, while CamCo [83] introduces epipolar constraints to improve geometric consistency. VD3D [1] embeds Plücker coordinates into transformer architectures for view synthesis, and ReCamMaster [4] focuses on re-rendering existing videos from novel camera paths. However, these methods focus on short-range, monocular sequences, limiting their applicability to large-scale multi-view environments in autonomous driving [24, 25, 75, 83, 1, 4]. OmniNWM extends camera controllability to panoramic, long-horizon settings through a normalized Plücker ray-map representation.
3 Methodology
At its core, OmniNWM unifies the state–action–reward triad within a single framework, enabling comprehensive state generation, precise action control, and integrated dense rewards. As illustrated in Figure 2, OmniNWM supports multi-modal generation (Section 3.1), producing pixel-aligned panoramic videos (Section 3.1.1), as well as 3D semantic occupancy (Section 3.1.2) and future planning trajectories (Section 3.1.3). Section 3.2 introduces the normalized panoramic Plücker ray-map camera control scheme, while Section 3.3 describes the flexible forcing strategy for long-term robust generation.
3.1 Comprehensive Generation within OmniNWM
The overview of the model architecture is illustrated in Figure 2. We leverage a pretrained 3D VAE [36] to encode input video frames into compact spatiotemporal latents with a compression ratio of $4\times 8\times 8$. A panoramic diffusion transformer (PDiT) with cross-view attention layers [79, 44] is leveraged to denoise the resulting latents. The denoised latents are decoded into pixel-aligned RGB, semantic, and depth outputs, which are subsequently used to produce 3D semantic occupancy and future planning trajectories.
3.1.1 Pixel-level Aligned Multi-model Generation
Joint generation of pixel-aligned multi-modal outputs has shown great potential to enhance the representational capacity of generative models and downstream performance [49, 43]. In OmniNWM, we initialize modality-specific noise latents for RGB, semantics, and depth, then concatenate them channel-wise for joint denoising using the PDiT. A shared decoder produces spatially aligned RGB, semantic, and depth video outputs. To ensure cross-modal consistency during training, semantic maps are colorized before VAE encoding and converted back to discrete labels via nearest-neighbor matching after decoding. The details of data curation on semantic and depth labels are provided in the supplementary material. This joint generation scheme ensures strong pixel-level alignment across modalities, which is essential for generating high-quality 3D occupancy, as detailed in the following subsection.
Figure 3: Occupancy generation. The generated panoramic RGB images, semantics, and depths are used to produce a 3D voxel volume. This design enables lightweight occupancy prediction compared with prior volumetric-based methods [43, 105], while naturally supporting the integration of occupancy-grounded rewards.
3.1.2 3D Occupancy Generation for Dense Rewards
As shown in Figure 3, the occupancy module generates a 3D voxel volume $\widehat{\mathbf{V}}$ from panoramic RGB, depth, and semantic inputs. Following previous work [44], a UNet based on a pre-trained EfficientNet-B7 [64] extracts image features $F_{i}$. Depth and semantic maps are processed via down-sampling and convolutional layers to produce depth features $F_{d}$ and semantic features $F_{s}$, respectively. SE3D blocks [30, 42] then perform adaptive aggregation between $F_{i}$ and $F_{d}$$F_{s}$. Finally, the 3D semantic voxel volume is computed by:
$\widehat{\mathbf{V}}=\texttt{Adap}_{d}(F_{i},F_{d})\otimes\texttt{Adap}_{s}(F_{i},F_{s}),$ (1)
where $\otimes$ denotes the outer product. $\texttt{Adap}_{d}$ and $\texttt{Adap}_{s}$ are adaptive aggregation networks for depth and semantic contexts. The resulting volume feature $\widehat{\mathbf{V}}$ is passed to a generation head with up-sampling and a softmax layer to produce the final 3D semantic occupancy. Unlike volumetric-based models like [43, 105] that directly synthesize full occupancy grids at high computational cost, our design computes occupancy from lightweight pixel-aligned features, which scales to long sequences and, crucially, enables integrated occupancy-grounded dense driving rewards.
Occupancy-grounded Rewards. The generated occupancy enables waypoint-wise reward computation following the class taxonomy in the NuScenes-Occupancy [71] dataset. The occupancy-grounded reward function $\widehat{\mathbf{R}}$ evaluates the safety, efficiency, and compliance of driving behaviors with detailed 3D environmental understanding as follows:
$\widehat{\mathbf{R}}=1+(R_{\text{col}}+R_{\text{bd}}+R_{\text{vel}})/N_{\text{reward}}.$ (2)
(1) The Collision Reward $R_{\text{col}}$ penalizes collisions between ego vehicle and obstacles, scaled by speed $v$ to reflect risks:
$R_{\text{col}}=-\alpha_{\text{col}}\cdot\mathbb{I}_{\text{col}}\cdot|v|,$ (3)
where $\mathbb{I}_{\text{col}}$ is a binary indicator function that equals 1 if a collision occurs and 0 otherwise. This is derived by querying the 3D occupancy for obstacle classes (e.g., ‘car’, ‘pedestrian’, ‘barrier’), and $\alpha_{\text{col}}$ is the balancing coefficient.
(2) The Boundary Reward $R_{\text{bd}}$ penalizes off-road behavior by detecting whether the ego vehicle leaves the drivable region (identified by the ‘drivable surface’ class):
$R_{\text{bd}}=-\alpha_{\text{bd}}\cdot\mathbb{I}_{\text{non-drivable}},$ (4)
where $\mathbb{I}_{\text{non-drivable}}$ is a binary indicator function that equals 1 when the vehicle is in a non-drivable area and 0 otherwise.
(3) The Velocity Reward $R_{\text{vel}}$ encourages maintaining a target speed of the ego vehicle to promote traffic efficiency:
$R_{\text{vel}}=-\alpha_{\text{vel}}\cdot\texttt{tanh}(|v-v_{\text{target}}|)\cdot\mathbb{I}_{v},$ (5)
where $v>0$, and $\mathbb{I}_{v}$ is a binary indicator function that equals 0 when $v\in[v_{min},v_{max}]$ and 1 otherwise.
By transforming 3D occupancy into dense waypoint-level feedback, the proposed reward function enables fine-grained evaluation of policy behavior under diverse scenarios. In a critical oncoming-truck scenario (Figure 4), the reward function effectively distinguishes driving quality: a high-speed collision (left) yields the lowest reward with severe $R_{\text{col}}$ penalty; an insufficient avoidance (middle) reduces reward; and a successful evasion (right) achieves the highest reward with minimal control penalty $R_{\text{col}}$ and velocity penalty $R_{\text{vel}}$, confirming discriminative capability. More evaluation results can be found in the supplementary.
Figure 4: Average rewards of different trajectories, computed using the proposed 3D occupancy-grounded reward function. The rewards effectively evaluate the feasibility of planning trajectories in the presence of obstacles (e.g., an oncoming truck).
3.1.3 Trajectory Planning for Closed-loop Navigation
As illustrated in Figure 1(c), OmniNWM functions as an environment simulator to enable closed-loop navigation with planning agents. Planning trajectories—generated from the model’s own outputs—guide future state generation when encoded into Plücker ray-maps.
To fully leverage OmniNWM’s multi-modal outputs, we developed a semantic-geometric reasoning vision-language-action (VLA) model based on Qwen-2.5-VL [3], termed OmniNWM-VLA, as the default planning agent. We introduce a plug-and-play Tri-Modal Mamba-based Interpreter (Tri-MIDI) to integrate semantic, depth, and RGB contexts before tokenization and input into OmniNWM-VLA. To adapt planning trajectories suitable as Plücker ray-map inputs, we extend the output channels of OmniNWM-VLA to predict not only the x and y coordinates of waypoints, but also their heading angles. In contrast to prior methods [11, 97], which produce key frames at 2HZ, our method operates at 12HZ on the NuScenes dataset. Note that for the initial video clip, the trajectory is produced by first generating multi-modal results (i.e., RGB, depth, and semantics) using a static trajectory, which are then fed into OmniNWM-VLA to enable multi-modal reasoning for the planner. Moreover, OmniNWM supports straightforward integration of various planning agents for closed-loop evaluation. Additional experimental comparisons are provided in Section 4.2, while further architectural and implementation details of OmniNWM-VLA are available in the supplementary.
3.2 Control with Panoramic Normalized Ray-map
Existing driving world models face two key limitations in action control: (1) sparse vectorized representations (e.g., waypoints) hinder accurate multi-view depiction [72, 102, 18], and (2) limited trajectory diversity in datasets restricts model generalization (Figure 5(c)). To overcome these issues, we propose converting trajectories into panoramic normalized Plücker ray-maps, providing a unified pixel-level representation that enables precise camera control (Figure 1(b)) and supports zero-shot generalization across datasets and camera configurations (Figure 10).
Figure 5: Panoramic normalized ray-map encoding. (a) Plücker ray-maps derived from panoramic camera poses are normalized in both scale and pose. (b) The ray-map normalization process constructs different trajectories within a unified 3D Plücker space. (c) Compared to the original NuScenes dataset, our strategy significantly enhances the diversity of trajectory distributions.
3.2.1 Efficient Plücker Ray-map Encoding
As illustrated in Figure 2, input trajectories are encoded into ray-maps via a parameter-free normalized Plücker encoder and injected into PDiT blocks. The normalized panoramic ray-maps are first downsampled in spatial and temporal dimensions to align with diffusion latents, and patchified into Plücker embedding tokens. These tokens are concatenated with the diffusion latent tokens (obtained via patchification of the diffusion latents) before being fed into the 3D full-attention layers within the PDiT blocks.
Following [24], Plücker ray-maps are constructed from camera intrinsics K{K} and extrinsics E=[R;t]{E}=[{R};{t}], where R{R} is the rotation matrix and t{t} the translation vector. For a pixel $(u,v)$, its Plücker embedding is defined as:
${p}_{u,v}=\left({o}\times{d}_{u,v},{d}_{u,v}\right)\in\mathbb{R}^{6},$ (6)
where o=t{o}={t} is the camera center in world coordinates, and ${d}_{u,v}$ is the direction vector from o{o} to the pixel, computed as:
${d}_{u,v}={R}{K}^{-1}[u,v,1]^{T}+{t}.$ (7)
The embedding of $i^{\text{th}}$ frame and $j^{\text{th}}$ view is represented as ${P}_{i,j}\in\mathbb{R}^{6\times h\times w}$, while a panoramic sequence of $n$ frames yields ${P}\in\mathbb{R}^{n\times 6\times v\times h\times w}$.
3.2.2 Scale and Pose Invariant Normalization
As shown in Figure 5 (a), we apply scale and pose normalization to each frame ${P}_{j}\in\mathbb{R}^{6\times 6\times h\times w}$ of the panoramic ray-map to obtain a unified Plücker representation. Our normalization method decouples camera parameters: a shared intrinsic matrix from a reference camera (defaulting to the front view) ensures scale consistency, while individual extrinsics retain distinct viewpoints. These parameters are then transformed into a unified coordinate system. For each pixel $(u,v)$ in the $k^{\text{th}}$ camera, the direction vector is computed using the reference intrinsics ${K}_{0}$ and the camera’s rotation matrix ${R}_{k}$:
$\hat{{d}}^{(k)}{u,v}={R}_{k}{K}_{0}^{-1}[u,v,1]^{T}.$ (8)
This projects all rays into a consistent metric scale. For pose invariance, both the camera center and direction vector are transformed into the reference camera’s coordinate system. The camera center becomes:
${o}_{k\to 0}={R}_{0}{R}_{k}^{T}({t}_{k}-{t}_{0}),$ (9)
and the direction vector is rotated accordingly:
$\hat{{d}}^{(k\to 0)}{u,v}={R}_{0}{R}_{k}^{T}\hat{{d}}^{(k)}_{u,v}$. (10)
The normalized Plücker coordinate is then given by:
$\hat{{p}}^{(k)}_{u,v}=\left({o}_{k\to 0}\times\hat{{d}}^{(k\to 0)}_{u,v},\hat{{d}}^{(k\to 0)}_{u,v}\right)$. (11)
This yields scale- and pose-normalized ray-maps $\hat{{P}}_{j}$ that are geometrically consistent across all views. This process constructs trajectories from different camera views within a unified 3D Plücker space (Figure 5(b)), significantly enriching the diversity of trajectory distributions (Figure 5(c)).
3.3 Long-term Generation with Flexible Forcing
To improve generation flexibility and robustness in long-term forecasting, we propose a flexible forcing strategy with multi-level noise injection. This approach corrupts preceding frames along multiple independent dimensions, enabling robust denoising under partially corrupted history. This is critical for mitigating error accumulation in extended auto-regressive generation [44, 14].
Figure 6: Flexible forcing strategy. (a) During training, independent multi-level noise is injected along view-wise and frame-wise dimensions. (b) During inference, flexible and robust generation is enabled via either frame-level auto-regression (many-to-one) or clip-level auto-regression (one-to-many or many-to-many).
Formally, during training, we independently inject noise into each frame $i$ and view $j$. Let $\mathbf{x}^{(i,j)}$ denote the latent representation for frame $i$ and view $j$. We corrupt it as:
$\tilde{\mathbf{x}}^{(i,j)}=\mathbf{x}^{(i,j)}+\alpha^{(i)}\cdot\epsilon_{\text{frame}}+\beta^{(j)}\cdot\epsilon_{\text{view}},$ (12)
where $\epsilon_{\text{frame}}\sim\mathcal{N}(0,I)$ and $\epsilon_{\text{view}}\sim\mathcal{N}(0,I)$ are independent noise vectors, and $\alpha^{(i)}$, $\beta^{(j)}$ are scaling factors. This per-view/frame noise scheme improves handling of spatio-temporal distortions compared to previous uniform noise application [43, 18].
As shown in Figure 6 (b), this formulation enables two distinct auto-regressive inference modes: (1) Frame-level auto-regression generates a single future frame conditioned on multiple preceding frames:
$\mathbf{\hat{x}}^{(t+1)}=f_{\theta}\left(\mathbf{\tilde{x}}^{(t-K)},\mathbf{\tilde{x}}^{(t-K+1)},\dots,\mathbf{\tilde{x}}^{(t)}\right),$ (13)
where $f_{\theta}$ is the diffusion model. $K$ is the input context window. $\mathbf{\tilde{x}}^{(t)}$ is the potentially corrupted latent at time $t$. This mode is particularly suitable for frame-wise trajectory planning simulation (Figure 4). (2) Clip-level auto-regression generates multiple future frames from fewer inputs:
$\mathbf{\hat{x}}^{(t+1)},\dots,\mathbf{\hat{x}}^{(t+L)}=f_{\theta}\left(\mathbf{\tilde{x}}^{(t-M)},\dots,\mathbf{\tilde{x}}^{(t)}\right),$ (14)
where $L$ is the output horizon and $M$ is the limited input context window. This approach is efficient for long-horizon generation while maintaining temporal coherence. During inference, historical frames can be partially noised to support both generation modes. For instance, in clip-level auto-regression, a single historical frame can be used as conditioning to generate multiple future frames initially, while multiple historical frames can be utilized in later stages to better exploit dynamic priors. Ablation studies in Figure 9 further validate the effectiveness of our approach.
4 Experiments
This section presents evaluation results of our approach against SOTA approaches across multiple applications.
4.1 Experimental Setup
We train our model on the NuScenes [6] and NuScenes-Occupancy [71] datasets. Our Panoramic Diffusion Transformer (PDiT) comprises 11.22B parameters, including an 11B Diffusion Transformer backbone consistent with prior works [109, 38] and 0.22B new parameters for cross-view attention layers. Training is conducted on 48 NVIDIA A800 GPUs with a batch size of 48. We employ the AdamW optimizer [50] with a learning rate of 1e-4. Further details regarding training objectives, experimental implementations of the occupancy generation module, and the OmniNWM-VLA are provided in the supplementary.
4.2 Main Results
Method Volumetric Condition-Free Multi-view Video FID $\downarrow$ FVD $\downarrow$
DriveGAN [35]
73.40 502.30
BEVControl [88]
24.85 -
DriveDreamer [72]
14.90 340.80
BEVGen [63]
25.54 -
DrivingGPT [10]
12.78 142.61
Vista [18]
6.90 89.40
DrivingWorld [31]
7.40 90.90
Epona [99]
7.50 82.80
WoVoGen [51]
27.60 417.70
X-Scene [91]
11.29 -
OccScene [44]
11.87 -
UniScene [43]
6.45 71.94
DiST-4D [21]
7.40 25.55
MagicDrive [17]
16.20 -
Panacea [78]
16.96 139.00
Drive-WM [73]
15.80 122.70
GenAD [106]
15.40 184.00
DriveDreamer-2 [102]
11.20 55.70
MagicDrive-V2 [16]
20.91 94.84
OmniNWM
5.45 23.63
Table 2: Quantitative evaluation of RGB video generation on the NuScenes validation set. Our proposed method outperforms previous SOTA methods in terms of generation quality.
Video Generation Quality. We evaluate RGB video generation quality using Frechet Inception Distance (FID) [26] and Frechet Video Distance (FVD) [67]. As shown in Table 2, our method achieves state-of-the-art performance of 5.45 FID and 23.63 FVD without requiring the heavy volumetric-based conditions (e.g., semantic occupancy [51, 44, 43], aggregated point clouds [21]), surpassing all previous approaches. These results demonstrate the effectiveness of our method for high-fidelity driving video generation.
Category Method Abs. Rel. $\downarrow$ $\delta<1.25$ $\uparrow$ $\delta<1.25^{2}$ $\uparrow$
Discriminative SurroundDepth [76] 0.28 0.66 0.84
M2Depth [112] 0.26 0.73 0.87
Generative Dist-4D [21] 0.39 0.58 0.81
OmniNWM 0.23 0.81 0.93
Table 3: Quantitative evaluation of generated panoramic depth on the NuScenes validation set.
Category Method Input
IoU
$\uparrow$
mIoU
$\uparrow$
[…]
LMSCNet [58] LiDAR 27.3 11.5
JS3C-Net [86] LiDAR 30.2 12.5
L-CONet [71] LiDAR 30.9 15.8
MonoScene [7] Camera 18.4 6.9
TPVFormer [32] Camera 15.3 7.8
C-CONet [71] Camera 20.1 12.8
HTCL [40] Camera 21.4 14.1
SparseOcc [65] Camera 21.8 14.1
Generative OccScene [44] Camera - 12.2
OccGen [68] Camera 23.4 14.5
OmniNWM Camera 33.3 19.8
Table 4: Quantitative results of generated semantic occupancy on the NuScenes-Occupancy validation set. Our method surpasses previous SOTA methods with a generative paradigm.
Figure 7: (a) Closed-loop evaluation of trajectory planning, reporting pass/fail counts, Scenario Pass Rate (SPR), and average rewards. (b) Distribution of average rewards across different methods on the NuScenes validation set.
Method Norm
RotErr
($10^{-2}$)
$\downarrow$
TransErr
($10^{-2}$)
$\downarrow$
GT Video - 1.36 1.75
UniScene [43] - 1.98 12.74
OmniNWM
2.11 14.19
OmniNWM
1.42 5.14
Table 5: Quantitative results of camera control accuracy on the NuScenes validation set. VGGT [69] is leveraged for camera pose extraction, and Similarity Transformation (Sim(3)) [28] is utilized for similarity scoring. Our method outperforms UniScene [43] obviously with ray-map normalization (‘Norm’).
Figure 8: Distribution of camera control performance with the rotation and translation errors on the NuScenes validation set.
We further evaluate the generated panoramic depth maps using standard metrics [48, 112]: absolute relative error (Abs. Rel.), and accuracy thresholds ($\delta$), with LiDAR-projected ground truth. As shown in Table 3, our method achieves superior performance on Abs. Rel. and $\delta$ metrics. Our method significantly outperforms the generative method of Dist-4D [21], and also surpasses the discriminative methods [76, 112], whose generalization ability is limited.
[…]
Trajectory Planning Evaluation. Figure 7 (a) presents closed-loop evaluation results for trajectory planning, showing Scenario Pass Rate (SPR) [61] and average rewards calculated across 150 scenes from the NuScenes validation set. Impromptu-VLA [11] achieves higher performance compared to Qwen-2.5-VL [3], with results comparable to our OmniNWM-VLA. Figure 7 (b) illustrates the distribution of average rewards—evaluated using our occupancy-grounded driving rewards—for different VLA planning baselines. The horizontal axis represents the percentage of the number of evaluation scenes. The results confirm that our reward function effectively distinguishes between the planning trajectories generated by these VLA planners.
Camera Control Accuracy. We evaluate camera control accuracy on the NuScenes validation set using Rotation Error (RotErr) and Translation Error (TransErr) metrics following previous works [24, 25]. Camera pose sequences are extracted using VGGT [69], with similarity scores calculated via Similarity Transformation (Sim(3)) [28]. As shown in Table 5 and Figure 8, our method with the panoramic normalized ray-map encoding significantly outperforms UniScene [43], yielding lower error distributions across the validation set. The RotErr achieved by our approach closely approaches ground-truth performance, while the higher TransErr reflects natural variability in driving distances across different scenarios. More evaluation results are provided in the supplementary.
RGB Semantics Depth IoU $\uparrow$ mIoU $\uparrow$
28.9 17.1
31.5 16.8
33.3 19.8
Table 6: Ablation study on the occupancy prediction module. The jointly generated semantics and depth effectively benefit occupancy prediction in contextual and geometric modeling.
Method FVD17 FVD33 FVD65 FVD129 FVD201
w/. flexible forcing 23.63 24.14 24.72 25.05 25.22
w/o. flexible forcing 26.79 59.64 102.79 249.74 386.72
Table 7: Ablation study on the flexible forcing strategy on different video lengths. Our strategy yields high-quality results with long-term generation sequences.
Zero-shot Generalization. Figure 10 presents zero-shot generalization results. We evaluate OmniNWM without fine-tuning across different datasets (e.g., nuplan or in-house collected datasets) and camera view configurations (e.g., 3 or 6 camera views). Leveraging the panoramic normalized ray-map encoding strategy with normalized Plücker ray-maps, our method demonstrates strong generalization capabilities across these diverse settings.
Figure 9: Visualization on the effect of flexible forcing strategy. Our approach effectively mitigates image degradation across multiple panoramic frames for long-term generation.
Figure 10: Zero-shot generalization across different new datasets and camera configurations without any fine-tuning.
4.3 Ablation Study
Occupancy Prediction. Table 6 presents ablation results for the occupancy prediction module. The baseline configuration predicts semantic occupancy from generated RGB, semantic, and depth maps within OmniNWM. The results indicate that the jointly generated semantic and depth maps yield noticeable improvements, enhancing the mIoU by 3.0 and 2.7, respectively. These results validate that the complementary benefits of the joint generated semantics and depth in contextual and geometric modeling.
Flexible Forcing. To validate the effectiveness of the flexible forcing strategy, we conduct ablation studies as shown in Table 7 and Figure 9. The results demonstrate that the proposed strategy effectively mitigates degradation for high-quality long-term generation, particularly in structural and textural details.
5 Conclusion and Future Work
OmniNWM establishes a new paradigm for driving world models by unifying state, action, and reward in a single omniscient framework. It jointly generates long-horizon, pixel-aligned panoramic videos of RGB, semantics, depth, and 3D occupancy (state); enables precise, generalizable camera control via normalized Plücker ray-maps (action); and introduces occupancy-grounded dense rewards for closed-loop evaluation (reward). This integration delivers state-of-the-art performance in generation fidelity, control accuracy, and zero-shot generalization—making OmniNWM a powerful and holistic framework for autonomous driving development. Limitations and Future Work. While highly capable, the model remains computationally intensive, motivating future work on efficient deployment.
Supplementary Material for OmniNWM
Appendix A More Related Works and Discussions
A.1 Semantic Occupancy Prediction
Semantic Occupancy Prediction (SOP), also referred to as Semantic Scene Completion (SSC), is a holistic 3D perception task that jointly infers the semantic labels and geometric completeness of a scene from partial observations [80, 39, 41, 13, 44, 100, 7, 32, 101, 74]. Existing methods can be broadly categorized into LiDAR-based and camera-based approaches. LiDAR-based methods benefit from precise 3D measurements and have advanced significantly through architectures designed to handle sparse and irregular point clouds. For instance, PolarNet [100] leverages a polar BEV representation to address uneven point distribution, enabling efficient real-time inference. SalsaNext [13] improves segmentation via residual dilated convolutions and Bayesian uncertainty estimation. RangeNet++ [54] uses range images and a custom post-processing step to mitigate discretization artifacts. SCPNet [80] introduces multi-scale feature aggregation and dense-to-sparse distillation to enhance representation learning. Camera-based methods have gained traction due to their lower cost and broader applicability. MonoScene [7] pioneers monocular 3D semantic completion by projecting 2D features into 3D space. StereoScene [39] further bridges stereo geometry and BEV features for improved semantic and geometric consistency. Subsequent works such as TPVFormer [32] introduce tri-perspective views for richer 3D representations, while SurroundOcc [77] and OccFormer [101] leverage multi-view images and transformer-based aggregation for dense occupancy estimation. Despite these advances, most existing methods rely on limited modalities, hindering their robustness in complex driving scenarios. In contrast, our approach first generates pixel-aligned multi-modal outputs—RGB, semantic, and metric depth—from a reference image, then leverages these complementary modalities to predict 3D semantic occupancy. This joint generation strategy fully exploits both geometric and contextual information, leading to more accurate and consistent 3D scene understanding.
A.2 Vision-Language Models
Recent years have witnessed growing interest in integrating Vision-Language Models (VLMs) and Large Language Models (LLMs) into autonomous driving systems, capitalizing on their strong reasoning capabilities and world knowledge [2, 104, 46, 59, 15, 85, 66, 107, 33, 108, 97]. These models often exhibit emergent reasoning skills, such as chain-of-thought (CoT) reasoning, which can be harnessed for interpretable decision-making. For example, DriveVLM [66] combines LLM-based trajectory proposals with an end-to-end refinement module for final planning. DriveGPT4 [85] uses LLMs in an iterative QA framework to explain driving behaviors and predict control signals. Doe-1 [107] reformulates driving as a next-token prediction task using a multi-modal generative model, while EMMA [33] encodes non-sensor inputs and outputs as natural language to fully leverage pre-trained LLMs. AutoVLA [111] utilizes raw visual inputs and language instructions for semantic reasoning and trajectory planning. FSDrive [97] introduces a pre-training paradigm for spatio-temporal reasoning, and Impromptu-VLA [11] provides a large-scale annotated dataset and baseline models for vision-language-action tasks in unstructured driving scenarios. In this work, we introduce OmniNWM-VLA, a semantic-geometric reasoning agent built upon Qwen-2.5-VL [2]. It integrates multi-modal inputs—RGB, depth, and semantic maps—via a Mamba-based [20] fusion block, enabling comprehensive scene understanding and trajectory planning.
Appendix B Architecture Overview of OmniNWM-VLA
As shown in Figure 11, OmniNWM-VLA is a semantic-geometric reasoning agent designed to bridge high-level scene understanding with precise trajectory planning in autonomous driving. Built upon Qwen-2.5-VL [2], it integrates multi-modal inputs—RGB, depth, and semantic maps—through a novel Mamba-based [20] fusion block, enabling comprehensive spatial reasoning and actionable trajectory prediction. The agent outputs not only waypoint coordinates but also heading angles, which are essential for generating normalized Plücker ray-maps used in panoramic video generation. In this way, our model supports precise control through Plücker ray-maps, bridging the gap between high-level reasoning and low-level action execution in a unified framework.
Figure 11: Architecture overview of OmniNWM-VLA. The Tri-Modal Mamba-based Interpreter (Tri-MIDI) fuses visual, geometric, and semantic information into a unified latent representation to forecast waypoint coordinates and heading angles.
Tri-Modal Mamba-based Interpreter. At the core of OmniNWM-VLA is the Tri-Modal Mamba-based Interpreter (Tri-MIDI), a lightweight and plug-and-play module that fuses visual, geometric, and semantic information into a unified latent representation. This design allows the model to reason jointly over appearance (RGB), 3D structure (depth), and scene semantics (segmentation), which is critical for forecasting geometrically consistent and semantically plausible trajectories.
Given aligned multi-view inputs—RGB images $X_{V}$, metric depth maps $X_{D}$, and semantic segmentation maps $X_{S}$—we first stitch them into a unified panoramic grid $P_{V},P_{D},P_{S}$ to maintain spatial consistency across views. Each modality is encoded separately using pre-trained encoders: CLIP [57] for RGB, SigLIP [98] for depth, and SegFormer [82] for semantics:
$H_{\alpha}=E_{\alpha}(P_{\alpha}),\quad\alpha\in\{V,D,S\}.$ (15)
The features are then projected into a common embedding space via modality-specific MLPs:
$Z_{\alpha}=\phi_{\alpha}(H_{\alpha}),\quad\alpha\in\{V,D,S\}.$ (16)
A Mamba-based state-space model $f_{LM}$ then performs cross-modal fusion, guided by a textual query $X_{T}$, to produce a fused latent representation:
$H_{R}=f_{LM}([Z_{V};Z_{D};Z_{S}],X_{T}).$ (17)
Using the Tokenized Rationale (TOR) mechanism [3], we insert special tokens into the query sequence to anchor intermediate reasoning steps. The output tokens at these positions are projected into the latent space of the downstream VLM:
$Z_{TOR}=\phi_{R}(H_{R_{tor}}).$ (18)
These tokens enrich the VLM’s input with spatially and semantically grounded cues, enabling it to predict trajectories that include both 2D waypoints and heading angles. This dual-output format is crucial for constructing normalized Plücker ray-maps, which provide pixel-level control over panoramic video generation in the OmniNWM framework.
By integrating multi-modal perception with sequential reasoning, OmniNWM-VLA supports closed-loop evaluation and long-horizon planning, effectively connecting high-level decision-making with low-level action execution in a unified autonomous driving pipeline.
Appendix C Semantic and Depth Data Curation
To construct semantic labels for the NuScenes dataset [6], we finetune a pretrained DDRNet model [27] on diverse driving scenes to enhance robustness. Training data includes the Cityscapes [12], Mapillary Vistas [56], Waymo Open [62], Woodscape [95], and BDD100k [96] datasets. Image-to-image translation techniques [34] are employed to synthesize night-scene imagery, further improving model generalization. The finetuned DDRNet generates semantic segmentation labels for the NuScenes dataset at 12 Hz.
For metric depth supervision, we follow established practices [21]: sparse LiDAR-projected depth and MVS-reconstructed depth maps, along with corresponding RGB images, are fed into a depth completion model [47] to generate dense, accurate depth maps.
Appendix D Training Objectives
Our framework comprises three core components: the Vision-Language-Action (VLA) planner, the Panoramic Diffusion Transformer (PDiT) for multi-modal video generation, and the 3D semantic occupancy prediction module. Each component is trained with specialized objectives tailored to its respective task.
VLA Planning Objective. The OmniNWM-VLA planner is built upon Qwen-2.5-VL [2] and inherits its training methodology. The model is optimized using standard causal language modeling objective, where the loss is computed as the negative log-likelihood of predicting the next token in the sequence:
$\mathcal{L}_{\text{VLA}}=-\mathbb{E}_{(x,y)\sim\mathcal{D}}\sum_{t=1}^{T}\log P(y_{t}|y_{<t},x),$ (19)
where $x$ represents the multi-modal input (concatenated RGB, depth, and semantic tokens processed by Tri-MIDI), $y$ is the target sequence containing both trajectory waypoints and heading angles, and $T$ is the sequence length. This objective enables the model to learn coherent multi-step reasoning for trajectory planning while maintaining compatibility with pre-trained VLMs.
Panoramic Video Generation Objective. For the PDiT-based panoramic video generation, we adopt the flow matching approach similar to OpenSora [109]. Let $X_{0}$ denote the video latent representation and $X_{1}\sim\mathcal{N}(0,1)$ represent Gaussian noise. The model $f_{\theta}$ takes as input an interpolated latent $X_{t}=(1-t)X_{0}+tX_{1}$ and is trained to predict the velocity component $X_{0}-X_{1}$. The corresponding loss function is formulated as:
$\mathcal{L}_{\text{PDiT}}=\mathbb{E}_{t,X_{0},X_{1}}\left[\left\|f_{\theta}(X_{t},t,y)-(X_{0}-X_{1})\right\|\right],$ (20)
where $y$ represents the conditioning input (historical trajectory and reference image), and the timestep $t$ is first sampled from a logit-normal distribution and then scaled according to the shape of $X_{0}$. This objective enables stable training and high-quality generation of panoramic RGB, semantic, and depth videos.
3D Semantic Occupancy Objective. For the 3D semantic occupancy prediction module, we follow the learning objective established in previous semantic scene completion works [7, 39]. The overall loss function combines multiple supervision signals:
$\mathcal{L}_{\text{Occ}}=\mathcal{L}_{\text{depth}}+\lambda_{\text{sem}}\mathcal{L}_{\text{sem}}+\lambda_{\text{geo}}\mathcal{L}_{\text{geo}}+\lambda_{\text{ce}}\mathcal{L}_{\text{ce}},$ (21)
where $\mathcal{L}_{\text{depth}}$ is the Binary cross-entropy loss that encourages accurate sparse depth distribution estimation. $\mathcal{L}_{\text{sem}}$ is the standard semantic loss for voxel-wise semantic classification. $\mathcal{L}_{\text{geo}}$ is the geometry loss that enforces structural consistency in the 3D volume. $\mathcal{L}_{\text{ce}}$ is the class-weighted cross-entropy loss to handle class imbalance. The $\lambda$ coefficients are balancing hyperparameters tuned to ensure stable training across different supervision signals. This multi-objective approach ensures that the generated occupancy volumes are both semantically accurate and geometrically plausible.
Appendix E Training Details
Panoramic Diffusion Transformer (PDiT). The training of our PDiT follows a progressive three-stage curriculum to ensure stable convergence and robust generalization:
Stage 1: Single-view Control (10K iterations). The model is initially trained on single-view sequences with Plücker ray-map control signals. We use a fixed frame length of 17 and an input resolution of $224\times 400$ to establish basic generation capabilities.
Stage 2: Multi-view and Multi-modal Extension (3K iterations). The model is extended to handle 6 panoramic camera views while incorporating joint generation of RGB, semantic, and depth modalities.
Stage 3: Variable-length and Resolution Training. Following [109], we introduce variability in sequence length (17 or 33 frames) and resolution ($224\times 400$ or $448\times 800$) to improve adaptability to diverse driving scenarios and computational constraints.
Throughout all stages, we employ the AdamW optimizer [50] with a learning rate of $1\times 10^{-4}$ and weight decay of 0.01. The model is trained on 48 NVIDIA A800 GPUs with a total batch size of 48.
3D Semantic Occupancy Objective. The occupancy generation module is implemented in PyTorch and trained for 24 epochs with a batch size of 8. We use the AdamW optimizer with a learning rate of $1\times 10^{-4}$ and weight decay of 0.01. Data augmentation includes random horizontal flipping and color jittering for RGB inputs.
Figure 12: Average rewards for specific scenarios across VLA planning baselines. Trajectories are evaluated using our occupancy-grounded reward function.
OmniNWM-VLA. Our OmniNWM-VLA planner follows the same training procedure as Qwen-2.5-VL [2], with modifications to accommodate autonomous driving requirements. The model uses a 3-B parameter architecture and is trained on a curated dataset of driving scenarios with multi-modal annotations. Unlike previous methods that operate at 2Hz, our model is trained at a control frequency of 12Hz to enable finer-grained trajectory planning. The training objective combines next-token prediction for trajectory waypoints and heading angles with multi-modal understanding tasks.
Overall Training Strategy. While each component can be trained independently, we employ a staged training strategy: first pre-training the PDiT and occupancy modules, then fine-tuning the VLA planner with the generated multi-modal outputs. This approach ensures that each module reaches optimal performance before being integrated into the full OmniNWM pipeline for closed-loop evaluation.
Appendix F More Visualization Results
Trajectory Planning Evaluation. Figure 12 presents the average rewards for specific scenarios on the NuScenes validation set across VLA planning baselines, evaluated using our occupancy-grounded driving reward. The results indicate that our reward function effectively distinguishes between the planning trajectories generated by different VLAs—such as those produced when the ego vehicle navigates near drivable area boundaries of median strips, or oncoming trucks.
Precise Panoramic Camera Control. As shown in the supplementary figures (Figures 13, 14, 15 and 16), including a case of reversing trajectories (Figure 17), OmniNWM enables precise panoramic camera control using a normalized Plücker ray-map representation. By converting diverse input trajectories into a unified pixel-level format, our method facilitates accurate viewpoint manipulation while preserving geometric consistency across camera views. This control mechanism is vital for generating realistic, physically plausible driving scenarios.
Long-term Navigation Stability. As shown in Figure 18, OmniNWM achieves robust long-term generation beyond ground-truth sequence lengths, producing coherent sequences of up to 321 frames. This capability is enabled by our flexible forcing strategy, which injects independent multi-level noise during training to mitigate error accumulation in auto-regressive inference. The results validate the model’s stability for extended closed-loop navigation simulations.
Comprehensive Multi-modal Generation. Figure 19 showcases OmniNWM’s joint generation of panoramic RGB, semantic segmentation, metric depth, and 3D occupancy outputs. The pixel-level alignment across modalities ensures geometric and semantic consistency, providing a comprehensive environment representation for downstream tasks such as trajectory planning and reward computation.
Figure 13: Precise panoramic camera control via normalized Plücker ray-maps. Given the same conditional frame, OmniNWM generates consistent multi-view videos by interpreting different input trajectories into pixel-level control signals.
Figure 14: Additional examples of panoramic camera control.
Figure 15: Additional examples of panoramic camera control.
Figure 16: Additional examples of panoramic camera control.
Figure 17: Precise panoramic camera control of reversing trajectories via normalized Plücker ray-maps.
Figure 18: Long-term navigation sequence (321 frames) generated through flexible forcing strategy. The model maintains temporal coherence and structural integrity beyond training sequence lengths, enabling extended closed-loop evaluation.
Figure 19: Joint multi-modal generation results, including panoramic RGB frames, semantic segmentation maps, metric depth estimations, and corresponding 3D semantic occupancy, demonstrating comprehensive scene generation capabilities.